Your firm is a U.K.-based exporter of shoes. You have sold an order to a French...

  1. Your firm is a U.K.-based exporter of shoes. You have sold an order to a French firm for €1,000,000 worth of shoes. Payment from the French firm (in euro) is due in 12 months. Use a money market hedge to redenominated this one-year receivable into a pound-denominated receivable with a one-year maturity.
    Contract Size Country U.S. $ equiv. Currency per U.S. $
    £ 10,000 Britain (pound) $ 1.9600 £ 0.5102 interest APR
    12 months forward $ 2.0000 £ 0.5000 rates
    10,000 Euro $ 1.5600 0.6410 i$ = 1 %
    12 months forward $ 1.6000 0.6250 i€ = 2 %
    SFr. 10,000 Swiss franc $ 0.9200 SFr. 1.0870 = 3 %
    12 months forward $ 1.0000 SFr. 1.0000 iSFr. = 4 %
    The following were computed without rounding. Select the answer closest to yours.

    £72,352.94

    £803,721.49

    €800,000

    £780,312.13
Answer #1

Receivable after one year = Euro 1,000,000

Borrow in Euro = 1,000,000/(1+2%) = Euro 980,392.16

Convert into Pound today = 980,392.16*1.56/1.96

= Pounds 780,312.1248

Invest for one year and get 780,312.13(1+3%) = Pound 803,721.49

Hence, the answer is

£803,721.49